OpenStudy (anonymous):

dy/dx for y= ln(7x^3-x^2)

OpenStudy (anonymous):

for ln(7x^3-2x^2) I believe you use the chain rule: f'(g)g' (1) so subtitute g in ln(g) g=7x^3-2x^2 (2) d/dx=21x^2-4x (3) f=ln(x): d/dx=1/x (4) using (1) above 1 --------- * 21x^2-4x 7x^3-2x^2 Now simplify etc 21x^2-4x --------- 7x^3-2x^2 (divide through by x) 21x-4 --------- 7x^2-2x